OpenStudy (compassionate):

F(x) = 12/(4x+2). Lets say we're finding the inverse f(x) = 12/(4x + 2) y = 12/(4x + 2) 4x + 2 = 12/y 4x + 2 * 12 = y 4x + 24 = y Now, we could just replace x and y 4y + 24 = x

OpenStudy (compassionate):

@campbell_st

OpenStudy (anonymous):

Are you asking a question? If you are finding the inverse; you should swap X and Y in the 2nd line...

OpenStudy (compassionate):

Hmm? Why?

OpenStudy (anonymous):

Maybe I should say it like this.. Swap F(x) with Y - as you did now solve for X then switch the variables

OpenStudy (anonymous):

Yes you did all of that and yes, just swap the variables as you stated. Sorry for the confusion

OpenStudy (compassionate):

F(x) = 12/(4x+2). y = 12/(4x + 2) 4x + 2 = 12/y 4x + 24 = y 4y + 24 = x

OpenStudy (compassionate):

@Hero , @Luigi0210

OpenStudy (anonymous):

Your line 3 is incorrect

OpenStudy (anonymous):

4x +2 = 12/y 4x = 12/y -2 x = 3/y - 1/2 now swap y = 3/x - 1/2

OpenStudy (compassionate):

Oh, I see :)

OpenStudy (anonymous):

You are welcome.